Course Home KHW 02 - Vectors Item 7 Express your answer with the appropriate units. Constants d 29.8 m John, Paul, and George are standing in a strawberry field. Paul is 12.0 m due west of John. George is 39.0mn from Paul, in a drection 34.0 south of east from Pauls location. Correct Part E What is the direction of Georges location from that of John? o south of east Submit XIncorrect; Try Again; 2 attempts remaining Return to Asignment Provide Feedback
Answer #1

Here ,

let east is the positive x direction and north is positive y direction

let the position of John is origin

Position of Paul , rp = -12 i m

position of George relative to Paul , rgp = 39 * (cos(34) i - sin(34)j)

Now, as rgp = rg - rp

rg = rp + rgp

rg = -12 i + 39 * (cos(34) i - sin(34)j)

rg = 20.33 i - 21.8 j

magnitude of displacement = sqrt(20.33^2 + 21.8^2) = 29.8 m
=============================================
part B)

direction of george relative to John = atan(21.8/20.33)

direction of george relative to John = 47 degree

the direction of george relative to John is 47 degree south of east